# Defining Neural Network Architecture through Polytope Structures of Dataset

### v1

@@ Appendix D Convergence on the proposed networks
In this section, we investigate whether gradient descent can converge to the networks that we proposed in the previous section. Specifically, we focus on the two-layer ReLU networks which are the basic building blocks of the constructions. Let $\mathcal{N}$ be a two-layer ReLU neural network defined in (1), where $\Theta:=\{v_{0}\}\cup\{v_{k},{\bm{w}}_{k},b_{k}\}_{k\in[0pt]}$ denotes the set of parameters of $\mathcal{N}$. For the given dataset ${{\mathcal{D}}}=\{({\bm{x}}_{i},y_{i})\}_{i=1}^{n}$, the mean squared error (MSE) loss and binary cross entropy (BCE) loss are defined by
@@
where $\ell(\mathcal{N},y):=\mathcal{N}y+(1-\mathcal{N})(1-y)$. Note that we introduce additional activation function $\sigma,\texttt{SIG}$ to define both loss functions.
We now employ a notion of ‘polyhedrally separable’ dataset from learning theory (Astorino & Gaudioso, 2002; Manwani & Sastry, 2010), which is a special case of polytope-basis cover; when given dataset can be separated by only one convex polytope as depicted in Figure 14 (a).
@@
We call that the dataset ${{\mathcal{D}}}$ is polyhedrally separable by $C$ if there exists a convex polytope $C$ such that ${\bm{x}}_{i}\in C$ if and only if $y_{i}=1$ for all $i\in[n]$.
We further introduce two notations. First, for a convex polytope $C$ composed of $0pt$ faces, we denote its $k$-th face by $\partial C_{k}$. Similarly, $\partial^{2}C_{k}$ denotes the boundary of $\partial C_{k}$, which refers to the ‘edge’ part of $C$. Second, for a set $A\subset{\mathbb{R}}^{d}$, #$\#(A):=|\{{\bm{x}}_{i}\in{{\mathcal{D}}}~{}|~{}{\bm{x}}_{i}\in A\}|$ denotes the number of data points ${\bm{x}}_{i}\in{{\mathcal{D}}}$ in $A$. We further need the following assumption, on the dataset ${{\mathcal{D}}}$ and network initialization.
Figure 14: Assumptions for the dataset and the network initialization. (a) Dataset ${{\mathcal{D}}}$ and a convex polytope $C$ satisfy the Assumption D.2. (b) One example of network initialization satisfying Assumption D.2. The red line displays the decision boundary ($\{{\bm{x}}\in{\mathbb{R}}^{d}~{}|~{}\mathcal{N}({\bm{x}})=0\}$).
Assumption D.2 (Dataset and initialization assumption).
@@
The dataset assumption (15) suggests that the data points in the set $\mathcal{B}_{r}(\partial C)$ for small $r$ are predominantly located in close proximity to the faces of the polytope $C$, rather than its corners (Figure 14(a)). The network initialization assumption implies that every neuron $({\bm{w}}_{k},b_{k})$ of $\mathcal{N}$ is initialized near $\partial C_{k}$ as described in Figure 14(b). With these assumptions, we can provide a discrete path that strictly decrease the loss value to zero.
@@
Suppose the dataset ${{\mathcal{D}}}$ and the two-layer network (1) satisfy Assumption D.2. Then,
@@
The proof of this theorem can be found in Appendix E.7. Theorem D.3 asserts that the loss landscape has no local minima in this initialization region. In other words, this theorem implies that stochastic (noisy) gradient descent is believed to converge to the global minimum, which has zero error on the training dataset ${{\mathcal{D}}}$. Therefore, the initialization conditions described in Assumption D.2, (17), and (18) can be understood as sufficient conditions such that gradient method converges to the global minimum.
@@ Appendix E Proofs
Figure 15: Proof of Proposition 3.1. (a) Given a hexagon which is approximated by $3$ hyperplanes given with blue lines and $1$ polytope in the second layer given with red hexagon. (b) Given a heptagon which is approximated by $4$ hyperplanes given with blue lines and $1$ polytope in the second layer given with red heptagon. (c) Hexagon has been extended to the $3$-dimensional polytope by incrementing the number of faces by $2$ with some potential first layer neurons (blue hyperplanes).
@@
We establish a terminology about the shape of prismatic polytopes. A prism in ${\mathbb{R}}^{3}$ consists of a ‘base’ and ‘height’ dimensions, and we generalize it to high dimensional prisms. We define a $k$-dimensional prismatic polytope in ${\mathbb{R}}^{d}$ as a topological space homeomorphic to $K\times{\mathbb{R}}^{d-k}$, where $\times$ denotes the Cartesian product an $K\subset{\mathbb{R}}^{k}$ is a compact set which is the ‘base’ of the prism. Roughly speaking, an $1$-dimensional prismatic polytope is a thick ‘hyperplane’ in ${\mathbb{R}}^{d}$, $(d-1)$-dimensional prismatic polytope is a long ’rod,’ and a $d$-dimensional prismatic polytope is just a convex polytope (Figure 16(a)). Then, for $k=1,2,\cdots,d$, removing a $k$-dimensional prism from $\mathcal{X}$ generates a $k$-dimensional hole, which increases the $(k-1)$-th Betti number $\beta_{k-1}$.
Now, we prove the theorem. Since $\mathcal{X}$ can be described as the union and difference of $\sum_{k=0}^{d-1}\beta_{k}$ convex polytopes, applying Theorem 3.4 provides
@@
For $1\leq k<d$, $\beta_{k}$ means the number of $k$-dimensional holes in $\mathcal{X}$, which was made by punching out a $k$-dimensional prismatic polytope. Since $k$-dimensional prisms have $2k$ faces that penetrate $\mathcal{X}$, we can reduce $2(d-k-1)$ number of hyperplanes that cover the hole. When $k=0$, it is easy to check that $2(\beta_{0}-1)$ hyperplanes are required to separate $\beta_{0}$ connected components. For example, Figure 16(c) shows this process for a topological space given in Figure 16(b). Then, the required total number of hyperplanes is bounded by
@@
Recall that $\mathcal{T}_{j}$ satisfies that $\sigma(\mathcal{T}_{j}({\bm{x}}_{i}))=0$ or 1 for all ${\bm{x}}_{i}\in{{\mathcal{D}}},j\in[J]$. From (5) and Lemma F.4, we know that $C_{j}:=\{{\bm{x}}~{}|~{}\mathcal{T}_{j}({\bm{x}})=1\}$ is a convex polytope for each $j\in[J]$. Then, we get
$\mathcal{N}({\bm{x}})$ $=\frac{1}{2}+\sum_{j=1}^{J}a_{j}\sigma(\mathcal{T}_{j}({\bm{x}}))$
$=\frac{1}{2}+\sum_{j=1}^{J}a_{j}\mathbbm{1}_{\{{\bm{x}}\in C_{j}\}}({\bm{x}}).$
@@
$\mathcal{N}({\bm{x}}_{i})>0\qquad\Leftrightarrow\qquad\sum_{C\in\mathcal{C}_{P}}\mathbbm{1}_{\{{\bm{x}}_{i}\in C\}}\geq\sum_{C\in\mathcal{C}_{Q}}\mathbbm{1}_{\{{\bm{x}}_{i}\in C\}}$
@@
Here we present proofs for each statement.
@@
    Secondly, we scrutinize the condition (7). Suppose there exists ${\bm{x}}_{i}\in{{\mathcal{D}}}$ such that $10<\mathcal{T}({\bm{x}}_{i})<$ 1. From Definition 5, it implies that
    $-1<\mathcal{T}({\bm{x}}_{i})-1=\sum_{k\in[m]}v_{k}\sigma({\bm{w}}_{k}^{\top}{\bm{x}}_{i}+b_{k})<0$
@@
    $\mathcal{T}({\bm{x}}_{i})-1\qquad\rightarrow\qquad\lambda_{scale}^{2}(\mathcal{T}({\bm{x}}_{i})-1).$
    Therefore, by repeating this scaling sufficiently many times, given $\lambda_{scale}>1$, $(\mathcal{T}({\bm{x}}_{i})-1)$ decreases beyon $-1$. This process applies to all such ${\bm{x}}_{i}$ in the dataset, eventually leading to $\sigma(\mathcal{T}({\bm{x}}_{i}))=0$ or 1 for all ${\bm{x}}_{i}\in{{\mathcal{D}}}$.
@@
    $C_{1}$ contains all points in ${{\mathcal{D}}}_{0}$. $C_{2}$ contains all points in ${{\mathcal{D}}}_{1}\cap C_{1}$. $C_{3}$ contains all points in ${{\mathcal{D}}}_{0}\cap C_{2}$. $\vdots$ $C_{J-1}$ contains all points in ${{\mathcal{D}}}_{\frac{1+(-1)^{J-1}}{2}}\cap C_{J-2}$. $C_{J}$ contains all points in ${{\mathcal{D}}}_{\frac{1+(-1)^{J}}{2}}\cap C_{J}$, and does not contain the another class.
@@
Note that we omitted $\Theta$ notation, the set of all learnable parameters. We will observe the change of these loss values with respect to one update of parameters. We add prime (${}^{\prime}$) for the updated parameter. For the given step size $\eta$, we explicitly provide the update of parameters by
@@ Appendix F Additional Propositions and Lemmas
Let $\mathcal{T}$ be a two-layer ReLU network defined in (5). Then, the classification region $R:=\{{\bm{x}}\in{\mathbb{R}}^{d}~{}|~{}\mathcal{T}({\bm{x}})>0\}$ is a convex polytope. Specifically, if the subset $S:=\{{\bm{x}}\in{\mathbb{R}}^{d}~{}|~{}\mathcal{T}({\bm{x}})=1\}$ is nonempty, then it is a convex polytope with $0pt$ faces.
@@
Now we consider the subset $S:=\{{\bm{x}}~{}|~{}\mathcal{T}({\bm{x}})=1\}$. Since all $v_{k}<0$, ${\bm{x}}\in S$ if and only if ${\bm{w}}_{k}^{\top}{\bm{x}}+b_{k}\leq 0$ for all $k\in[0pt]$. Then, $S$ is a convex polytope with $0pt$ faces by Definition 2.1. ∎

### v2

@@ Appendix D Convergence on the Proposed Networks
In this section, we investigate whether gradient descent can converge to the networks we proposed in the main text (cf. Theorem 3.4). Specifically, we focus on two-layer ReLU networks, which are the basic building blocks of the constructions. Let $\mathcal{N}$ be a two-layer ReLU neural network defined in (1), where $\Theta:=\{v_{0}\}\cup\{v_{k},{\bm{w}}_{k},b_{k}\}_{k\in[0pt]}$ denotes the set of parameters of $\mathcal{N}$. For the given dataset ${{\mathcal{D}}}=\{({\bm{x}}_{i},y_{i})\}_{i=1}^{n}$, we consider a binary classification problem under the following two loss functions: the mean squared error (MSE) loss and binary cross entropy (BCE) loss. They are defined by
@@
where $\ell(\mathcal{N},y):=\mathcal{N}y+(1-\mathcal{N})(1-y)$. Note that we introduce additional activation function $\sigma$ and SIG to define both loss functions. Specifically, we adopt additional ReLU activation on the output layer for the existence of the zero-loss solution in (13).
We now employ the notion of ‘polyhedrally separable’ dataset from the learning theory (Astorino & Gaudioso, 2002; Manwani & Sastry, 2010), which is a special case of polytope-basis cover; when a given dataset can be separated by only one convex polytope as depicted in Figure 15 (a).
@@
We say that the dataset ${{\mathcal{D}}}=\{({\bm{x}}_{i},y_{i})\}_{i\in[n]}$ is polyhedrally separable by $C$ if there exists a convex polytope $C$ such that ${\bm{x}}_{i}\in C$ if and only if $y_{i}=1$ for all $i\in[n]$.
We further introduce two notations. First, for a convex polytope $C$ composed of $0pt$ faces, we denote its $k$-th face by $\partial C_{k}$. Similarly, $\partial^{2}C_{k}$ denotes the boundary of $\partial C_{k}$, which refers to the ‘edge’ part of $C$. Second, for a set $A\subset{\mathbb{R}}^{d}$, #$\#(A):=|\{{\bm{x}}_{i}\in{{\mathcal{D}}}~{}|~{}{\bm{x}}_{i}\in A\}|$ denotes the number of data points ${\bm{x}}_{i}\in{{\mathcal{D}}}$ in $A$. We further need the following assumptions on the dataset ${{\mathcal{D}}}$ and network initialization.
Figure 15: Assumptions for the dataset and the network initialization. (a) Dataset ${{\mathcal{D}}}$ and a convex polytope $C$ satisfy the Assumption D.2. (b) One example of network initialization satisfying Assumption D.2. The red line displays the decision boundary of $\mathcal{N}$.
Assumption D.2 (Dataset and initialization assumptions).
@@
The dataset assumption (15) suggests that the data points in the set $\mathcal{B}_{r}(\partial C)$ for small $r$ are predominantly located in close proximity to the faces of the polytope $C$, rather than its corners (Figure 15(a)). The network initialization assumption implies that every neuron $({\bm{w}}_{k},b_{k})$ of $\mathcal{N}$ is initialized near $\partial C_{k}$ as described in Figure 15(b). With these assumptions, we can prove the existence of a discrete path that strictly decrease the loss value to zero.
@@
Suppose the dataset ${{\mathcal{D}}}$ and the two-layer network $\mathcal{N}$ in (1) satisfy Assumption D.2. Then,
@@
The proof of this theorem can be found in Appendix E.7. Theorem D.3 asserts that the loss landscape has no local minima within this initialization region. If local minima did exist in this region, it would contradict the presence of a loss-decreasing path from the local minima to the global minima. Consequently, this theorem provides strong evidence for the convergence of gradient descent to the global minima.
However, it is important to note that there may still be saddle points where gradient descent could potentially get stuck. In such cases, we believe that stochastic (noisy) gradient descent may help in escaping these saddle points and eventually converging to the global minimum, which has zero error on the training dataset ${{\mathcal{D}}}$. Therefore, the initialization conditions described in Assumption D.2, (17), and (18) can be understood as necessary conditions for ensuring that the gradient method converges to the global minimum.
Lastly, we mention that Theorem D.3 can be easily extended to the three-layer network (7) proposed in Theorem 3.4. For such a three-layer network $\mathcal{N}$, Theorem D.3 can be applied to each two-layer subnetwork $\mathcal{T}_{j}$ to generate the loss-decreasing path. By combining all these paths, a unified loss-decreasing path for $\mathcal{N}$ is formed. This extension underscores the robustness and generality of the convergence properties demonstrated, ensuring that even more complex network architectures retain the desirable characteristics of gradient descent convergence.
@@ Appendix E Proofs
Figure 16: Proof of Proposition 3.1.
@@
Remark E.1.
Rigorously, the lower bound on the network width proposed in Proposition 3.1 can also be understood as the maximum number of faces that a given network can approximate with its polytope. Conversely, to achieve the UAP and approximate any polytope with $0pt$ faces, the width of the first hidden layer must be greater than or equal to $0pt$. This is precisely explained in Proposition F.6, which proves that the upper bound proposed in Proposition 3.1 is tight and sufficient to satisfy the UAP.
@@
We establish a terminology about the shape of prismatic polytopes. A prism in ${\mathbb{R}}^{3}$ consists of a ‘base’ and ‘height’ dimensions, and we generalize it to high dimensional prisms. We define a $k$-dimensional prismatic polytope in ${\mathbb{R}}^{d}$ as a topological space homeomorphic to $K\times{\mathbb{R}}^{d-k}$, where $\times$ denotes the Cartesian product an $K\subset{\mathbb{R}}^{k}$ is a compact set which is the ‘base’ of the prism. A bounded $k$-dimensional prismatic polytope is a intersection of a $k$-dimensional prismatic polytope and a bounded convex polytope.
Roughly speaking, an $1$-dimensional prismatic polytope is a thick ‘hyperplane’ in ${\mathbb{R}}^{d}$, $(d-1)$-dimensional prismatic polytope is a long ’rod,’ and a $d$-dimensional prismatic polytope is just a convex polytope, as shown in Figure 17(a). Then, for $k=1,2,\cdots,d$, removing a $k$-dimensional prism from $\mathcal{X}$ generates a $k$-dimensional hole, which increases the $(k-1)$-th Betti number $\beta_{k-1}$. In Figure 17(b), we provide an example of such cover $\mathcal{X}$ in ${\mathbb{R}}^{3}$. $\mathcal{X}$ is described by subtracting six prismatic polytopes from a large bunoid in ${\mathbb{R}}^{3}$. In this case, the subtracted prismatic polytopes can be understood as a bounded prismatic polytopes with six faces.
Now, we prove the theorem. Recall that the polytope-basis cover $\mathcal{X}$ can be described as a subtraction of $\sum_{k=0}^{d-1}\beta_{k}$ convex polytopes from a sufficiently large convex polytope with $m$ faces. Applying Theorem 3.4, we get
@@
This upper bound of $d_{1}$ can be further reduced. For $1\leq k<d$, $\beta_{k}$ means the number of $k$-dimensional holes in $\mathcal{X}$, which was made by punching out a $k$-dimensional prismatic polytope. Since $k$-dimensional prisms have $2k$ faces that penetrate $\mathcal{X}$, we can reduce $2(d-k-1)$ number of hyperplanes that cover the hole. When $k=0$, it is easy to check that $2(\beta_{0}-1)$ hyperplanes are required to separate $\beta_{0}$ connected components. For instance, Figure 17(c) shows this process for a topological space given in Figure 17(b). Then, the required total number of hyperplanes is bounded by
@@
Recall that $\mathcal{T}_{j}$ satisfies that $\sigma(\mathcal{T}_{j}({\bm{x}}_{i}))=0$ or $\lambda$ for all ${\bm{x}}_{i}\in{{\mathcal{D}}},j\in[J]$. From (5) and Lemma F.4, we know that $C_{j}:=\{{\bm{x}}~{}|~{}\mathcal{T}_{j}({\bm{x}})=\lambda\}$ is a convex polytope for each $j\in[J]$. Then, we get
$\mathcal{N}({\bm{x}})$ $=-\frac{1}{2}\lambda+\sum_{j=1}^{J}a_{j}\sigma(\mathcal{T}_{j}({\bm{x}}))$
$=-\frac{1}{2}\lambda+\sum_{j=1}^{J}a_{j}\mathbbm{1}_{\{{\bm{x}}\in C_{j}\}}({\bm{x}}).$
@@
$\mathcal{N}({\bm{x}}_{i})>0\qquad\Longleftrightarrow\qquad\sum_{C\in\mathcal{C}_{P}}\mathbbm{1}_{\{{\bm{x}}_{i}\in C\}}>\sum_{C\in\mathcal{C}_{Q}}\mathbbm{1}_{\{{\bm{x}}_{i}\in C\}}$
@@
Here, we present proofs for each statement.
@@
    Secondly, we scrutinize the condition (7). Suppose there exists ${\bm{x}}_{i}\in{{\mathcal{D}}}$ such that $0<\mathcal{T}({\bm{x}}_{i})<\lambda$. From Definition 5, it implies that
    $-\lambda<\mathcal{T}({\bm{x}}_{i})-\lambda=\sum_{k\in[m]}v_{k}\sigma({\bm{w}}_{k}^{\top}{\bm{x}}_{i}+b_{k})<0$
@@
    $\mathcal{T}({\bm{x}}_{i})-\lambda\qquad\rightarrow\qquad\lambda_{scale}^{2}(\mathcal{T}({\bm{x}}_{i})-\lambda).$
    Therefore, by repeating this scaling sufficiently many times, given $\lambda_{scale}>1$, $(\mathcal{T}({\bm{x}}_{i})-\lambda)$ decreases under $-1$. This process applies to all such ${\bm{x}}_{i}$ in the dataset, eventually leading to $\sigma(\mathcal{T}({\bm{x}}_{i}))=0$ or $\lambda$ for all ${\bm{x}}_{i}\in{{\mathcal{D}}}$.
@@
    $C_{1}$ contains all points in ${{\mathcal{D}}}_{0}$.
    $C_{2}$ contains all points in ${{\mathcal{D}}}_{1}\cap C_{1}$.
    $C_{3}$ contains all points in ${{\mathcal{D}}}_{0}\cap C_{2}$.
    $\vdots$
    $C_{J-1}$ contains all points in ${{\mathcal{D}}}_{\frac{1+(-1)^{J-1}}{2}}\cap C_{J-2}$.
    $C_{J}$ contains all points in ${{\mathcal{D}}}_{\frac{1+(-1)^{J}}{2}}\cap C_{J}$, and does not contain the another class.
@@
Note that we omitted $\Theta$ notation, the set of all learnable parameters. We will observe the change of these loss values with respect to one update of parameters. We add prime () for the updated parameter. For the given step size $\eta$, we explicitly provide the update of parameters by
@@ Appendix F Additional Propositions and Lemmas
Let $\mathcal{T}$ be a two-layer ReLU network defined in (5). Then, the classification region $R:=\{{\bm{x}}\in{\mathbb{R}}^{d}~{}|~{}\mathcal{T}({\bm{x}})>0\}$ is a convex polytope. Specifically, if the subset $S:=\{{\bm{x}}\in{\mathbb{R}}^{d}~{}|~{}\mathcal{T}({\bm{x}})=\lambda\}$ is nonempty, then it is a convex polytope with $0pt$ faces.
@@
Now we consider the subset $S:=\{{\bm{x}}~{}|~{}\mathcal{T}({\bm{x}})=\lambda\}$. Since all $v_{k}<0$, ${\bm{x}}\in S$ if and only if ${\bm{w}}_{k}^{\top}{\bm{x}}+b_{k}\leq 0$ for all $k\in[0pt]$. Then, $S$ is a convex polytope with $0pt$ faces by Definition 2.1. ∎
@@
Proposition F.6.
Consider the neural network architecture $d\stackrel{{\scriptstyle\sigma}}{{\rightarrow}}d_{1}\stackrel{{\scriptstyle\sigma}}{{\rightarrow}}d_{2}\stackrel{{\scriptstyle\sigma}}{{\rightarrow}}\cdots\stackrel{{\scriptstyle\sigma}}{{\rightarrow}}d_{D}\rightarrow 1$ and a convex polytope $C$ with $m$ faces. Then,
  1. 1.
    if $d_{1}\geq m$ and $d_{2}\geq 1$, then it is a feasible architecture on $C$.
  2. 2.
    if $\max_{j}d_{j}\leq m-1$, then it may not be a feasible architecture on some polytope $C$.
  3. 3.
    if $d_{1}\leq m-2$, then it may not be a feasible architecture on some polytope $C$.
Proof.
The proof is accomplished by two strategies: for a feasible architecture, we explicitly construct such neural networks. For the negative statements, we prove them by providing some counterexamples.
  1. 1.
    Proposition 3.1 shows that $d\stackrel{{\scriptstyle\sigma}}{{\rightarrow}}m\rightarrow 1$ is a feasible architecture . Therefore, for $d_{1}\geq m$, then taking the identity for all other layers, it becomes a feasible architecture .
  2. 2.
    When $d_{1}\leq 0pt-1$, there is a $0pt$-faces convex polytope $C$ that cannot be approximated by the given network architecture. The simplest example is a half-space ($10pt=$ 1).
    Below, we provide another non-trivial example: Let $C$ be a $d$-simplex in ${\mathbb{R}}^{d}$, thus $10pt=d+$ 1. Suppose $\max_{j}d_{j}\leq 0pt-1=d$. Then, by Lemma F.7, we conclude that the classified regions are always unbounded. Therefore, it cannot approximate a bounded polytope $C$.
  3. 3.
    From the above proof, recall the $d$-simplex $C$ (thus $m=d+1$). If $d_{1}\leq m-2=d-1$, we provide a counterexample proving that it cannot be approximated by a ReLU network with the architecture $d\stackrel{{\scriptstyle\sigma}}{{\rightarrow}}(m-2)\stackrel{{\scriptstyle\sigma}}{{\rightarrow}}\cdots\stackrel{{\scriptstyle\sigma}}{{\rightarrow}}d_{D}\rightarrow 1$. Let ${\bm{w}}_{1},\cdots,{\bm{w}}_{0pt-2}$ be the weight vectors of the first layer. Since the dimension of $span<{\bm{w}}_{k}>$ is equal or less than $10pt-2=d-$ 1, there exists a nonzero vector $\hat{\bm{w}}\in span<\{{\bm{w}}_{k}\}_{k\in[m]}>^{\perp}$. In other words, $\hat{\bm{w}}^{\top}{\bm{w}}_{k}=0$ for all $k\in[m-2]$. Then, it implies $\mathcal{N}({\bm{x}}+t\hat{\bm{w}})=\mathcal{N}({\bm{x}})$ for all $t\in{\mathbb{R}}$. Therefore, $\mathcal{N}$ cannot approximate the bounded polytope $C$.
Lemma F.7.
Let $\mathcal{N}$ be a deep ReLU network where all hidden dimension is equal or smaller than the input dimension $d$. Suppose $\mu(\{\mathcal{N}({\bm{x}})>0\})>0$. Then, $\mu(\{\mathcal{N}({\bm{x}})>0\})$ is either 00 or $\infty$. In other words, the classification region is either measure-zero or unbounded.
Proof.
Beise et al. (2021, Theorem 2) showed that if all hidden layers have width equal or smaller than the input dimension, then the connected components of every decision region are unbounded. Therefore, $\mu(\{\mathcal{N}({\bm{x}})>0\})$ is either 00 or $\infty$, depends on whether $\{\mathcal{N}({\bm{x}})>0\}$ is empty or not. ∎